OpenStudy (anonymous):

If the integral from 2 to 4 of f(x)dx=6, then the integral from 2 to 4 of (f(x)+3)dx= Answer: 12 Please explain how to come to this answer.

hartnn (hartnn):

\(\Large \int \limits_2^4[f(x)+3 ]dx = \int \limits_2^4f(x) dx +\int \limits_2^43 dx\)

hartnn (hartnn):

you already know 1st integral, can you try to find 2nd ?

OpenStudy (anonymous):

Oh that makes sense. Thank you for your help! :)

hartnn (hartnn):

most welcome ^_^